# FLAIR-HUB: Large-scale Multimodal Dataset for Land Cover and Crop Mapping Anatol Garioud, Sébastien Giordano, Nicolas David, Nicolas Gonthier Institut national de l’information géographique et forestière (IGN), France

### v1

@@ V Baseline architecture
TRAIN VAL TRAIN$+$VAL TEST ALL
@@ B.1 Land Cover Mapping
Model Input Data Supervision Pixel ×$\times 10^{9}$ PARA. FLAIR #1 FLAIR #2
@@ B.2 Crop Type Mapping
Limitations of Modality Contribution: Configurations that combine multi-temporal inputs (S1, S2, or both) with high-resolution imagery (aerial VHR or SPOT) generally achieve the best performance for the LPIS task. This likely results from the complementarity of modalities: high-resolution imagery provides fine-grained textural features, while time series data capture phenological dynamics critical for distinguishing between crop types. However, performance differences across specific modality combinations are not always intuitive. For instance, the Aerial+S2 configuration (LPIS-C) does not consistently outperform either of the single-modality baselines such as Aerial (LPIS-A) or S2 (LPIS-F). In some cases, the differences are substantial: for example, barley and rapeseed exhibit IoU drops of 21% and 12.2%, respectively, in LPIS-C compared to LPIS-F. A similar pattern is observed when comparing LPIS-I and LPIS-J: adding aerial imagery to a configuration using SPOT and Sentinel-1/2 time series (LPIS-I) results in large IoU decreases for Flax/Hemp ($-40.6$%) and Soy ($-18.1$%). These findings suggest that simply adding modalities does not guarantee improved performance. The observed inconsistencies may stem from limitations in both the dataset and the model architecture. Potential contributing factors include increased model complexity leading to overfitting, interactions between the main and auxiliary losses when adding modalities, and limitations of the current temporal encoder. These factors indicate a need for more effective fusion strategies and potentially stronger architectures to fully leverage multimodal data.

### v2

@@ V Baseline architecture
TRAIN VAL TRAIN++VAL TEST ALL
@@ B.1 Land Cover Mapping
Model Input Data Supervision Pixel $\times 10^{9}$ PARA. FLAIR #1 FLAIR #2
@@ B.2 Crop Type Mapping
Limitations of Modality Contribution: Configurations that combine multi-temporal inputs (S1, S2, or both) with high-resolution imagery (aerial VHR or SPOT) generally achieve the best performance for the LPIS task. This likely results from the complementarity of modalities: high-resolution imagery provides fine-grained textural features, while time series data capture phenological dynamics critical for distinguishing between crop types. However, performance differences across specific modality combinations are not always intuitive. For instance, the Aerial+S2 configuration (LPIS-C) does not consistently outperform either of the single-modality baselines such as Aerial (LPIS-A) or S2 (LPIS-F). In some cases, the differences are substantial: for example, barley and rapeseed exhibit IoU drops of 21% and 12.2%, respectively, in LPIS-C compared to LPIS-F. A similar pattern is observed when comparing LPIS-I and LPIS-J: adding aerial imagery to a configuration using SPOT and Sentinel-1/2 time series (LPIS-I) results in large IoU decreases for Flax/Hemp (40.6-40.6%) and Soy (18.1-18.1%). These findings suggest that simply adding modalities does not guarantee improved performance. The observed inconsistencies may stem from limitations in both the dataset and the model architecture. Potential contributing factors include increased model complexity leading to overfitting, interactions between the main and auxiliary losses when adding modalities, and limitations of the current temporal encoder. These factors indicate a need for more effective fusion strategies and potentially stronger architectures to fully leverage multimodal data.